Solve sin’(x) + sin(x) = 0 for 0 SI<2m. Give your answer in radians.
Answer #1

SOLUTION;

Find out the solution of

Sin^{2}x+sinx=0

Where 0\leq x< 2\pi

Now,

Sin^{2}x+sinx=0

Sinx(sinx+1)=0

so,

sinx+1=0. And.   sinx=0

sin x= -1

sinx=-1

x=\frac{3\pi}{2}

and

Sinx=0

x=0,\pi

the answer is

x=0,\pi,\frac{3\pi}{2}